OpenStudy (anonymous):

Easy Math Basics Question 9x · 9(4) · 4x · -4(4)

OpenStudy (anonymous):

So i'll post how far i'v got now 9x · 36 · 4x · (-16)

OpenStudy (anonymous):

Now my problem is trying to understand how like terms work, would 9x and 4x combine into 13x or \[36^2\]?

OpenStudy (anonymous):

correct so far, now just rearrange the x's and the numbers (commutative property of multipication) and multiply.

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

alright i'll look up commutative property i'm only familiar with distributive so far thanks

OpenStudy (anonymous):

okay I stil don't understand this haha

OpenStudy (anonymous):

when i multiply 9x into 4x is my output 13x or 36x^2?

OpenStudy (anonymous):

The calculator gives me 36x^2 but i don't know if that is correct for this scenario

OpenStudy (alekos):

yes it is

OpenStudy (anonymous):

ah okay thanks! haha commutative property led me to a · b = b · a which didn't help much

OpenStudy (alekos):

9 times 4 = 36 and x times x = x^2 then you just go 36 times x^2 = 36x^2

OpenStudy (anonymous):

ya i thought 9x · 4x = 13x i'm silly